### v2

@@ Abstract
Adaptive inference schemes reduce the cost of machine learning inference by assigning smaller models to easier examples, attempting to avoid invocation of larger models when possible. In this work we explore a simple, effective adaptive inference technique we term Agreement-Based Cascading (ABC). ABC builds a cascade of models of increasing size/complexity, and uses agreement between ensembles of models at each level of the cascade as a basis for data-dependent routing. Although ensemble execution introduces additional expense, we show that these costs can be easily offset in practice due to large expected differences in model sizes, parallel inference execution capabilities, and accuracy benefits of ensembling. We examine ABC theoretically and empirically in terms of these parameters, showing that the approach can reliably act as a drop-in replacement for existing models and surpass the best single model it aims to replace in terms of both efficiency and accuracy. Additionally, we explore the performance of ABC relative to existing cascading methods in three common scenarios: (1) edge-to-cloud inference, where ABC reduces communication costs by up to 14×14$\times$14 ×; (2) cloud-based model-serving, where it achieves a 3×3$\times$3 × reduction in rental costs; and (3) inference via model API services, where ABC achieves a $2$-25×25$\times$25 × reduction in average price per token/request relative to state-of-the-art LLM cascades.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
The high cost of inference associated with deploying large machine learning (ML) models presents a significant barrier to their adoption (Strubell et al., 2020; Kaplan et al., 2020). As models continue to increase in size, practitioners are often faced with investing substantial resources in updating existing deployments or settling for lower-performing alternatives. However, for many applications, it has been shown that a considerable portion of the data seen during inference can be effectively evaluated using small models rather than large, state-of-the-art models Chen et al. (2020); Jitkrittum et al. (2023). This means that if we can identify the subset of data samples that can be accurately evaluated by more inexpensive models, average inference costs can be reduced. This problem is often referred to as adaptive inference, where the cost of inference adapts to some notion of ‘difficulty’ of each example seen at inference time.
@@
A natural approach for adaptive inference is to cascade over a set of potential models, starting from the least expensive and moving to more expensive models based on some deferral rule Rowley et al. (1998); Viola & Jones (2001); Soo (2014). A common cascade construction is to use a Pareto-efficient set of models and an easy-to-compute deferral rule such as the confidence scores of the models predictions Viola & Jones (2004); Wang et al. (2018a; 2021). Recently, cascading has seen renewed interest in the ML community with the advent of large language models (LLMs) and vision foundation models (e.g., Chen et al., 2023; Gupta et al., 2024). Recent approaches consider a variety of techniques to construct model cascades, including designing novel model architectures that have built-in cascading capabilities (Cai et al., 2019; Devvrit et al., 2023; Khare et al., 2023), or learning routing schemes/deferral rules that require data-dependent training for every task considered (Chen et al., 2023; Ding et al., 2024). While these approaches can improve accuracy/efficiency trade-offs, they may also introduce significant computational overhead in setup and training/finetuning costs.
@@
Figure 1: (a) Agreement-Based Cascading (ABC): ABC introduces a data-dependent routing scheme that uses agreement amongst an ensemble of models to determine whether to cascade to larger models. If the predictions of smaller ensembles do not align, the cascade moves to the next tier of larger models, continuing until agreement is reached or the largest model(s) are used. This can reduce cost by limiting the use of the largest models to cases where smaller models cannot reach consensus. (b) ABC is a natural baseline to consider for adaptive inference due to (i) the vast number of pretrained models available to ML practitioners today; (ii) the fact that even small accuracy gains often require an order-of-magnitude increase in FLOPs, mirroring proposed scaling laws and resulting in large differences in model sizes between cascade tiers Hestness et al. (2017); Henighan et al. (2020); Madaan et al. (2023). The pink dashed line represents the Pareto-optimal frontier, showing the models with the highest accuracy for a given computational budget. We show that ABC can effectively improve this frontier—allowing practitioners to achieve high accuracy without incurring the full computational cost of the largest models by selectively invoking smaller models for “easy” samples.
@@
We point to a few trends in ML that make ABC a particularly attractive approach for model cascading. In particular, while using an ensemble of models at each level may initially appear to increase overall inference costs, it is natural to believe this simple baseline could excel in real-world applications due to: (1) the growing ease of obtaining pretrained models of various sizes and accuracies due to the rise of model registries Wolf et al. (2019) as well as flexible compression schemes Zhu et al. (2023); Dennis et al. (2023); (2) scaling laws that predict a large increase in inference cost for every marginal increase in accuracy  Hestness et al. (2017) (Figure 1); and (3) an increased ability to execute ensembles of relatively small models in parallel, with minimal additional costs Fern & Givan (2003); Kim et al. (2023); Miao et al. (2023) (Figure 3).
@@
  • We propose ABC as a training-free, deferral-based cascading approach that leverages agreement among ensembles of existing models, eliminating the need for additional routing networks or specialized architectures.
@@ 2 Related Work
Adaptive inference schemes have been a topic of interest in machine learning for many years. This section discusses three main approaches to cascading and adaptive inference: score-based methods, trained routing procedures, and dynamic networks. We highlight how ABC’s approach relates to and differs from these existing methods.
@@ 2.1 Cascading using Score-based Deferrals
Traditional cascading methods often rely on simple score-based metrics for deferral decisions. We compare to the recent Wisdom-of-Committees method of Wang et al. (2021) as a general, representative method in this category but note that specialized instantiations of this approach have been applied across various domains in prior work, for instance, in object detection (Rowley et al., 1998; Viola & Jones, 2004; Wang et al., 2011; Cai et al., 2015; Angelova et al., 2015; Streeter, 2018), image classification (Wang et al., 2018a; 2021), and text classification (Li et al., 2021b; Mamou et al., 2022; Varshney & Baral, 2022; Lebovitz et al., 2023). In these settings, deferrals within cascading systems use metrics such as confidence scores, entropy, and probabilities (Gangrade et al., 2021; Geifman & El-Yaniv, 2019; Narasimhan et al., 2022). While using confidence scores of existing models is inexpensive, these scores are required to be well-calibrated, which is less common for off-the-shelf models (Guo et al., 2017; Enomoro & Eda, 2021). Recent methods have thus considered a variety of techniques to enhance score-based deferrals, such as explicitly adding deferrals as a prediction (Wang et al., 2018a), difficulty-aware regularization (Li et al., 2021b), temperature scaling Wang et al. (2023), and calibration mechanisms (Nie et al., 2024).
@@ 2.2 Cascades with Routing Procedures
Other methods avoid confidence scores entirely by training procedures independent of the model set to route instances as needed. This trend has gained prominence with the rise of black-box model API services for LLMs, where prediction scores are often unavailable. Guan et al. (2018) developed a selection module trained to determine the best-fit classifiers for data instances. Yue et al. (2024) (MoT LLM Cascade) used sampling and consistency checking to determine when to defer to a high-cost model. AutoMix (Madaan et al., 2023) proposed a few-shot self-verification mechanism, similar to Yue et al. (2024), but also introduced a Markov-based meta-verifier for cascading in context-grounded tasks. FrugalGPT (Chen et al., 2023) leveraged a cascade strategy that triages incoming queries using a DistilBERT router and scoring function. HybridLLM Ding et al. (2024) used a fine-tuned DeBERTa He et al. (2020) to route queries to models based on the predicted query difficulty and the desired quality level. On the same note, OrchestraLLM Lee et al. (2023)—using hand-labeled data to create expert model pools—selects model to query based on embedding distances between the pools’ instances and the test instances. Other methods also used some form of trained router, including RouteLLM Ong et al. (2024), ‘Fly-swat or cannon’ Šakota et al. (2024), and Shnitzer et al. (2023).
@@ 2.3 Dynamic and Adaptive Networks
Routing procedures typically train an independent routing mechanism, keeping the Pareto-set of models untouched. However, for many medium-scale applications, methods have been explored that learn a Pareto-set of models and stopping criterion together. For example, early exit methods of (Huang et al., 2018; Wang et al., 2018b; Xin et al., 2020; Geng et al., 2021; Zhou et al., 2020; Liu et al., 2020; Schuster et al., 2022), subnetworks extraction methods, (Yu et al., 2018; Yu & Huang, 2019; Chen et al., 2021; Hou et al., 2020; Han et al., 2022; Devvrit et al., 2023), composition-based methods (Suggala et al., 2020; Dennis et al., 2023; Du & Kaelbling, 2024) are some of the more recent examples.
@@ 3 Agreement-Based Cascading (ABC)
In this section, we present a high-level overview of the Agreement-Based Cascading (ABC) approach, providing the essential concepts and insights needed to understand ABC ’s effectiveness in the experiments (Section 5). Readers interested in the formalizations and technical details can refer to Section 4 for a more comprehensive treatment.
@@ 3.1 Overview of ABC Approach
  Current cascade level, $i\leftarrow 1$
  Cascaded prediction, $y\leftarrow\emptyset$
  for $i\in{\{1,\dots,n_{E}\}}$ do
     $y\leftarrow H_{i}(x)$
     if $r_{i}(x)=0$ then
        break // Models in ensemble ‘agree’
     end if
  end for
  return $y$
@@ 3.2 ABC in Modern ML Environments
In recent years, obtaining trained models has become much easier, thanks to the rise of public and private model repositories Wolf et al. (2019), where a wide variety of pre-trained models are readily available. This abundance of accessible models of various sizes and performance levels makes the practical application of ABC especially appealing. Unlike methods that require additional task-specific training or fine-tuning, ABC can leverage these existing models, making it suitable for direct deployment as a “drop-in” replacement for high-cost models. This adaptability, combined with minimal setup and training requirements, can allow ABC to fit seamlessly into many existing ML workflows.
Strictly speaking, evaluating agreement between multiple models in an ensemble is expensive when compared to the small router models that are used in many existing approaches. However, two key aspects of modern ML workloads can help to mitigate this additional cost in practice. First, in many cases some degree of parallelization is available that can reduce the impact on inference cost metrics such as latency. We use $\rho$ to smoothly interpolate between the fully sequential case $(\rho=0)$ to the fully parallel case, $(\rho=1)$ as detailed in Section 4.3. Second, in many use-cases the difference in cost between models in successive tiers of cascades is so large that the impact of lower tier models on the overall cost of inference is negligible, even with the added cost of constructing ensembles (see Figure 3). We use $\gamma$ to denote the relative cost of the models — the ratio of the cost of the smaller model to the larger model.
In Section 5, we evaluate ABC ’s performance across several real-world scenarios (resulting in varying settings of $\rho$ and $\gamma$), demonstrating its competitiveness and efficiency. We first show that ABC’s ensemble-based deferral mechanism preserves and often improves on the accuracy of models it aims to replace. We also show that when the lower tier models are small enough compared to the higher tier models, ABC ’s ensemble-based deferral rule operates with negligible impact on the overall cost.
@@ 4 Formalization and Theoretical Details
In this section, we formalize our problem setup by first outlining the standard statistical learning framework, which serves to define the concepts of models, ensembles, the learning performance (eg. accuracy) and the inference cost associated with ensembles (Section 4.1). Next, we describe deferral rule-based cascades and formulate the specific case of drop-in cascades, cascades which prioritise accuracy over inference cost savings (Section 4.2). We introduce safe deferral rules as a sufficiency condition for constructing drop-in cascades and define our deferral rule based on agreement between models in an ensemble (Section 4.3). We conclude this section by formalizing the test-time accuracy-inference cost behaviour of drop-in cascades (Section 4.4).
@@ 4.1 Problem Setup and Notation
As a concrete example, in the classification setup, where we use the miss-classification error as a loss function, the risk is given by $\mathcal{R}(h)=\mathbb{E}(y\neq h(x))$.
@@ 4.2 Deferral Rule-Based Drop-in Cascade
In this work, we focus on deployment scenarios where a decrease in model performance is a critical concern. This is particularly relevant when cascades are intended to function as a drop-in replacement for an existing deployment of an expensive model. This notion can be formulated as maximizing the number of calls to the smaller model, while retaining the accuracy of the larger model. With slight abuse of notation, let $\mathcal{M}_{r}(x)$ denote the prediction of the cascade with =$\mathcal{M}={\{H^{k}_{1},h_{2}\}}$ and deferral rule $r$. Then we desire that for a choice of small error budget $\xi>0$,
@@
The constant function $r(x)=1$, which defers for every $x$, is feasible and attains the objective value of 00. Moreover, every feasible deferral rule leads to a cascade with competitive accuracy as $h_{2}$. Instead of picking an error budget $\xi$, a complimentary (dual) approach is to consider a fixed inference budget and aim to attain the best accuracy within the budget. In this view, drops in accuracy (compared to the existing large model) are permissible, provided the inference budget is met.  Jitkrittum et al. (2023) examine this perspective, which we recommend to interested readers.

### v3

@@ Abstract
Adaptive inference schemes reduce the cost of machine learning inference by assigning smaller models to easier examples, attempting to avoid invocation of larger models when possible. In this work we explore a simple, effective adaptive inference technique we term Agreement-Based Cascading (ABC). ABC builds a cascade of models of increasing size/complexity and uses agreement between ensembles of models at each level of the cascade as a basis for data-dependent routing. Although ensemble execution introduces additional expense, we show that these costs can be easily offset in practice due to large expected differences in model sizes, parallel inference execution capabilities, and accuracy benefits of ensembling. We examine ABC theoretically and empirically in terms of these parameters, showing that the approach can reliably act as a drop-in replacement for existing models and surpass the best single model it aims to replace in terms of both efficiency and accuracy. Additionally, we explore the performance of ABC relative to existing cascading methods in three common scenarios: (1) edge-to-cloud inference, where ABC reduces communication costs by up to 14×14$\times$14 ×; (2) cloud-based model serving, where it achieves a 3×3$\times$3 × reduction in rental costs; and (3) inference via model API services, where ABC achieves a 222-25×25$\times$25 × reduction in average price per token/request relative to state-of-the-art LLM cascades.
@@ 1 Introduction
The high cost of inference associated with deploying large machine learning (ML) models presents a significant barrier to their adoption (Strubell et al., 2020; Kaplan et al., 2020). As models continue to increase in size, practitioners are often faced with investing substantial resources in updating existing deployments or settling for lower-performing alternatives. However, for many applications, it has been shown that a considerable portion of the data seen during inference can be effectively evaluated using small models rather than large, state-of-the-art models (Chen et al., 2020; Jitkrittum et al., 2023). This means that if we can identify the subset of data samples that can be accurately evaluated by more inexpensive models, average inference costs can be reduced. This problem is often referred to as adaptive inference, where the cost of inference adapts to some notion of ‘difficulty’ of each example seen at inference time.
@@
A natural approach for adaptive inference is to cascade over a set of potential models, starting from the least expensive and moving to more expensive models based on some deferral rule (Rowley et al., 1998; Viola & Jones, 2001; Soo, 2014). A common cascade construction is to use a Pareto-efficient set of models and an easy-to-compute deferral rule such as the confidence scores of the models predictions (Viola & Jones, 2004; Wang et al., 2018a; 2021). Recently, cascading has seen renewed interest in the ML community with the advent of large language models (LLMs) and vision foundation models (e.g., Chen et al., 2023; Gupta et al., 2024). Recent approaches consider a variety of techniques to construct model cascades, including designing novel model architectures that have built-in cascading capabilities (Cai et al., 2019; Devvrit et al., 2023; Khare et al., 2023), or learning routing schemes/deferral rules that require data-dependent training for every task considered (Chen et al., 2023; Ding et al., 2024). While these approaches can improve accuracy/efficiency trade-offs, they may also introduce significant computational overhead in setup and training/finetuning costs.
@@
Figure 1: (a) Agreement-Based Cascading (ABC): ABC introduces a data-dependent routing scheme that uses agreement amongst an ensemble of models to determine whether to cascade to larger models. If the predictions of smaller ensembles do not align, the cascade moves to the next tier of larger models, continuing until agreement is reached or the largest model(s) are used. This can reduce cost by limiting the use of the largest models to cases where smaller models cannot reach consensus. (b) ABC is a natural baseline for adaptive inference due to (i) the vast number of pretrained models available to ML practitioners today; (ii) the fact that even small accuracy gains often require an order-of-magnitude increase in FLOPs, mirroring proposed scaling laws and resulting in large differences in model sizes between cascade tiers (Hestness et al., 2017; Henighan et al., 2020; Madaan et al., 2023). The pink dashed line represents the Pareto-optimal frontier, showing the models with the highest accuracy for a given computational budget. We show that ABC can effectively improve this frontier—allowing practitioners to achieve high accuracy without incurring the full computational cost of the largest models by invoking smaller models for ‘easier’ samples.
@@
We point to a few trends in ML that make ABC a particularly attractive approach for model cascading. In particular, while using an ensemble of models at each level may initially appear to increase overall inference costs, it is natural to believe this simple baseline could excel in real-world applications due to: (1) the growing ease of obtaining pretrained models of various sizes and accuracies due to the rise of model registries (Wolf et al., 2019) as well as flexible compression schemes (Zhu et al., 2023; Dennis et al., 2023); (2) scaling laws that predict a large increase in inference cost for every marginal increase in accuracy  (Hestness et al., 2017) (Figure 1); and (3) an increased ability to execute ensembles of relatively small models in parallel, with minimal additional costs (Fern & Givan, 2003; Kim et al., 2023; Miao et al., 2023) (Figure 3).
@@
  • We propose ABC as a training-free, deferral-based cascading approach that leverages agreement among ensembles of existing models, removing the need for additional routing networks or specialized architectures.
@@ 2 Related Work
Adaptive inference schemes have been a topic of interest in machine learning for many years. This section discusses three main approaches to cascading and adaptive inference: score-based methods, trained routers, and dynamic networks. We highlight how ABC’s approach relates to and differs from these existing methods.
@@ 2.1 Cascading using Score-based Deferrals
Traditional cascading methods often rely on simple score-based metrics for deferral decisions. We compare to the recent Wisdom-of-Committees method of Wang et al. (2021) as a general, representative method in this category but note that specialized instantiations of this approach have been applied across various domains in prior work, for instance, in object detection (Rowley et al., 1998; Viola & Jones, 2004; Wang et al., 2011; Cai et al., 2015; Angelova et al., 2015; Streeter, 2018), image classification (Wang et al., 2018a; 2021), and text classification (Li et al., 2021b; Mamou et al., 2022; Varshney & Baral, 2022; Lebovitz et al., 2023). In these settings, deferrals within cascading systems use metrics such as confidence scores, entropy, and probabilities (Gangrade et al., 2021; Geifman & El-Yaniv, 2019; Narasimhan et al., 2022). While using confidence scores of existing models is inexpensive, these scores are required to be well-calibrated, which is less common for off-the-shelf models (Guo et al., 2017; Enomoro & Eda, 2021). Recent methods have thus considered a variety of techniques to enhance score-based deferrals, such as explicitly adding deferrals as a prediction (Wang et al., 2018a), difficulty-aware regularization (Li et al., 2021b), temperature scaling (Wang et al., 2023), and calibration mechanisms (Nie et al., 2024).
@@ 2.2 Cascades with Routing Procedures
Other methods avoid confidence scores entirely by training procedures independent of the model set to route instances as needed. This trend has gained prominence with the rise of black-box model API services for LLMs, where prediction scores are often unavailable. Guan et al. (2018) developed a selection module trained to determine the best-fit classifiers for data instances. Yue et al. (2024) (MoT LLM Cascade) used sampling and consistency checking to determine when to defer to a high-cost model. AutoMix (Madaan et al., 2023) proposed a few-shot self-verification mechanism, similar to Yue et al. (2024), but also introduced a Markov-based meta-verifier for cascading in context-grounded tasks. FrugalGPT (Chen et al., 2023) leveraged a cascade strategy that triages incoming queries using a DistilBERT router and scoring function. HybridLLM (Ding et al., 2024) used a fine-tuned DeBERTa (He et al., 2020) to route queries to models based on the predicted query difficulty and the desired quality level. On the same note, OrchestraLLM (Lee et al., 2023)—using hand-labeled data to create expert model pools—selects model to query based on embedding distances between the pools’ instances and the test instances. Other methods also used some form of trained router, including RouteLLM (Ong et al., 2024), ‘Fly-swat or cannon’ (Šakota et al., 2024), and Shnitzer et al. (2023).
@@ 2.3 Dynamic and Adaptive Networks
Routing procedures typically train an independent routing mechanism, keeping the Pareto-set of models untouched. However, for many medium-scale applications, methods have been explored that learn a Pareto-set of models and stopping criterion together. For example, early exit methods of Bolukbasi et al. (2017); Huang et al. (2018); Shafiee et al. (2018); Wang et al. (2018b); Hu et al. (2020); Xin et al. (2020); Geng et al. (2021); Zhou et al. (2020); Liu et al. (2020); Schuster et al. (2022), subnetworks extraction methods (Yu et al., 2018; Yu & Huang, 2019; Chen et al., 2021; Hou et al., 2020; Han et al., 2022; Devvrit et al., 2023), composition-based methods (Suggala et al., 2020; Dennis et al., 2023; Du & Kaelbling, 2024) are some of the more recent examples. While these methods achieve impressive efficiency gains, they require training specialized architectures from scratch with adaptive capabilities.
@@ 3 Agreement-Based Cascading (ABC)
In this section, we present a high-level overview of the Agreement-Based Cascading (ABC) approach, providing the essential concepts and insights needed to understand ABC ’s effectiveness in the experiments (§5). Readers interested in the formalizations and technical details can refer to §4 for a more comprehensive treatment.
@@ 3.1 Overview of ABC Approach
1: Current cascade level, $i\leftarrow 1$
2: Cascaded prediction, $y\leftarrow\emptyset$
3:for $i\in{\{1,\dots,n_{E}\}}$ do
4:  $y\leftarrow H_{i}(x)$
5:  if $r_{i}(x)=0$ then
6:   break {Models in ensemble ‘agree’}
7:  end if
8:end for
9:return $y$
@@ 3.2 ABC in Modern ML Environments
In recent years, obtaining trained models has become much easier, thanks to the rise of public and private model repositories (Wolf et al., 2019), where a wide variety of pre-trained models are readily available. This abundance of accessible models of various sizes and performance levels makes the practical application of ABC especially appealing. Unlike methods that require additional task-specific training or fine-tuning, ABC can leverage these existing models, making it suitable for direct deployment as a “drop-in” replacement for high-cost models. This adaptability, combined with minimal setup and training requirements, can allow ABC to fit seamlessly into many existing ML workflows.
Strictly speaking, evaluating agreement between multiple models in an ensemble is expensive when compared to the small router models that are used in many existing approaches. However, two key aspects of modern ML workloads can help to mitigate this additional cost in practice. First, in many cases some degree of parallelization is available that can reduce the impact on inference cost metrics such as latency. We use $\rho$ to smoothly interpolate between the fully sequential case $(\rho=0)$ to the fully parallel case, $(\rho=1)$ as detailed in §4.3. Second, in many use-cases the difference in cost between models in successive tiers of cascades is so large that the impact of lower tier models on the overall cost of inference is negligible, even with the added cost of constructing ensembles (see Figure 3). We use $\gamma$ to denote the relative cost of the models — the ratio of the cost of the smaller model to the larger model.
In §5, we evaluate ABC ’s performance across several real-world scenarios (resulting in varying settings of $\rho$ and $\gamma$), demonstrating its competitiveness and efficiency. We first show that ABC’s ensemble-based deferral mechanism preserves and often improves on the accuracy of models it aims to replace. We also show that when the lower tier models are small enough compared to the higher tier models, ABC ’s ensemble-based deferral rule operates with negligible impact on the overall cost.
@@ 4 Formalization and Theoretical Details
In this section, we formalize our problem setup by first outlining the standard statistical learning framework, which serves to define the concepts of models, ensembles, the learning performance (e.g., accuracy) and the inference cost associated with ensembles (§4.1). Next, we describe deferral rule-based cascades and formulate the specific case of drop-in cascades, cascades which prioritise accuracy over inference cost savings (§4.2). We introduce safe deferral rules as a sufficiency condition for constructing drop-in cascades and define our deferral rule based on agreement between models in an ensemble (§4.3). We conclude this section by formalizing the test-time accuracy-inference cost behaviour of drop-in cascades (§4.4).
@@ 4.1 Problem Setup and Notation
As a concrete example, in the classification setup, where we use the mis-classification error as a loss function, the risk is given by $\mathcal{R}(h)=\mathbb{E}(y\neq h(x))$.
@@ 4.2 Deferral Rule-Based Drop-in Cascade
In this work, we focus on deployment scenarios where a decrease in model performance is a critical concern. This is particularly relevant when cascades are intended to function as a drop-in replacement for an existing deployment of an expensive model. This notion can be formulated as maximizing the number of calls to the smaller model while retaining the accuracy of the larger model. With a slight abuse of notation, let $\mathcal{M}_{r}(x)$ denote the prediction of the cascade with =$\mathcal{M}={\{H^{k}_{1},h_{2}\}}$ and deferral rule $r$. Then we desire that for a choice of a small error budget $\xi>0$,
@@
The constant function $r(x)=1$, which defers for every $x$, is feasible and attains the objective value of 0. Moreover, every feasible deferral rule leads to a cascade with competitive accuracy as $h_{2}$. Instead of picking an error budget $\xi$, a complimentary (dual) approach is to consider a fixed inference budget and aim to attain the best accuracy within the budget. In this view, drops in accuracy (compared to the existing large model) are permissible, provided the inference budget is met.  Jitkrittum et al. (2023) examine this perspective, which we recommend to interested readers.